In 20X2, sales increased to $5,740,000 and the assets for that year were as follows:

cash $163,000

Accounts receivable 924,000

inventory 1,063,000

net plant and equipment 520,000

total assets $2,670,000

Compute the following: round your answers to 2 decimal places (times)

Accounts receivable turnover

inventory turnover

fixed asset turnover

total asset turnover
Answer #1

accounts receivable turnover = sales/account receivable

= 5740000/924000

= 6.21

inventory turnover = sales/inventory

= 5740000/1063000

= 5.41

Fixed asset turnover = sales/net plant and equipment

= 5740000/520000

= 11.04

total asset turnover = sales/total assets

= 5740000/2670000

= 2.16